Castle, Inc., has no debt outstanding and a total market value of $240,000. Earnings before interest and taxes, EBIT, are projected to be $36,000 if economic conditions are normal. If there is strong expansion in the economy, then EBIT will be 20 percent higher. If there is a recession, then EBIT will be 25 percent lower. The firm is considering a debt issue of $155,000 with an interest rate of 6 percent. The proceeds will be used to repurchase shares of stock. There are currently 6,000 shares outstanding. Ignore taxes for questions a and b. Assume the stock price remains constant.

a-1. Calculate return on equity (ROE) under each of the three economic scenarios before any debt is issued. (Do not round intermediate calculations. Enter your answers as a percent rounded to 2 decimal places, e.g., 32.16.)
  

ROE
Recession %
Normal %
Expansion %


a-2.
Calculate the percentage changes in ROE when the economy expands or enters a recession. (A negative answer should be indicated by a minus sign. Do not round intermediate calculations. Enter your answers as a percent rounded to the nearest whole number, e.g., 32.)

  

% change in ROE
Recession %
Expansion %


Assume the firm goes through with the proposed recapitalization.

b-1.
Calculate the return on equity (ROE) under each of the three economic scenarios. (Do not round intermediate calculations. Enter your answers as a percent rounded to 2 decimal places, e.g., 32.16.)

  

ROE
Recession %
Normal %
Expansion %


b-2.
Calculate the percentage changes in ROE when the economy expands or enters a recession. (A negative answer should be indicated by a minus sign. Do not round intermediate calculations. Enter your answers as a percent rounded to 2 decimal places, e.g., 32.16.)

  

% change in ROE
Recession %
Expansion %


Assume the firm has a tax rate of 35 percent.
Answer #1

a-1) Using the following Data: debt issue EBIT (Normal $240,000 $36,000 change in case of expansion % change in case of recesb-1) Market Value before Debt Issue Debt Issued Market Value after debt Issue 35 $240,000 $155,000 $85,000 36 38 39 InterestFormula sheet

A B C D E F G H I J
2 a-1)
3 Using the following Data:
4 Market Value of Equity before debt issue 240000
5 EBIT (Normal) 36000
6 % change in case of expansion 0.2
7 % change in case of recession -0.25
8 Tax rate 0
9 Using the above data, since debt and taxes are zero, therefore net income can be calculated as following:
10
11 EBIT Interest Expense Income Before Tax Tax Expense Net Income
12 Recession =D13*(1+D7) 0 =D12-E12 =F12*$D$8 =F12-G12 =F12-G12
13 Normal =D5 0 =D13-E13 =F13*$D$8 =F13-G13
14 Expansion =D5*(1+D6) 0 =D14-E14 =F14*$D$8 =F14-G14
15
16 Return on Equity calculation:
17 Return on Equity is given by following formula:
18
19
20
21
22 Since preferred dividend is zero, Return on Equity can be calculated as follows:
23
24 Net Income Shareholder's Equity Return on Equity
25 Recession =H12 =$D$4 =D25/E25 =D25/E25
26 Normal =H13 =$D$4 =D26/E26
27 Expansion =H14 =$D$4 =D27/E27
28
29 a-2)
30
31 % change in ROE
32 Recession =(F25-F26)/F26 =(F25-F26)/F26
33 Expansion =(F27-F26)/F26
34
35 b-1)
36 Market Value before Debt Issue 240000
37 Debt Issued 155000
38 Market Value after debt Issue =D36-D37
39
40 Interest Rate 0.06
41 Interest Expense =D37*D40
42 Tax Rate 0
43
44 Using the above data, net income can be calculated as following:
45
46 EBIT Interest Expense Income Before Tax Tax Expense Net Income
47 Recession =D12 =$D$41 =D47-E47 =F47*$D$42 =F47-G47 =F47-G47
48 Normal =D13 =$D$41 =D48-E48 =F48*$D$42 =F48-G48
49 Expansion =D14 =$D$41 =D49-E49 =F49*$D$42 =F49-G49
50
51 Return on Equity calculation:
52 Return on Equity is given by following formula:
53
54
55
56
57 Since preferred dividend is zero, Return on Equity can be calculated as follows:
58
59 Net Income Shareholder's Equity Return on Equity
60 Recession =H47 =$D$38 =D60/E60 =D60/E60
61 Normal =H48 =$D$38 =D61/E61
62 Expansion =H49 =$D$38 =D62/E62
63
64 b-2)
65
66 % change in ROE
67 Recession =(F60-F61)/F61 =(F60-F61)/F61
68 Expansion =(F62-F61)/F61
69